Assume that in an economy with taxes and imports, the MPC = 0.75, t=0.25 and m = 0.3. The multiplier is
1
Expert's answer
2019-04-15T10:31:59-0400

If

MPC = 0.75   - marginal propensity to consume

t=0.25          - taxes

m = 0.3.        - the limiting inclination to import

Then

government spending multiplier:

GSM= 1/(1-MPC)=1/(1-0.75)=4

government spending multiplier with taxes:

GSMt= 1/(1-MPC(1-t))=1/(1-0.75*0.7)=2.1

government spending multiplier in an open economy (with tax and import):

GSMoe=1/(1-MPC(1-t)+m)=1/(1-0.75*0.7+0.3)=1.29